Variable costs for Sheridan Company are 20% of sales. Its selling price is $125 per unit. If Sheridan sells one unit more than break-even units, how much will profit increase?

$100

$625

$25

$50

Please try to explain how to do the calculations for the answers! thanksssss
Answer #1

Answer: $100

Variable cost Per Unit = 125 *20% = 25

Contribution MArgin = 125-25 =$100

At the Break-Even Point

total Cost = Total Revenue

After Break-even Point the Profit of the Firm Will increase by Contribution Margin Per Unit for Each Unit of sale.